the chairs for a concert are arranged so that each row has the same number of chairs. sarah has 4 chairs in front of her, and 8 chairs behind her. there are 5 seats to her left and 6 to her right. how many chairs are set up?
Answer by edjones(8007)

4+1+8=13 rows
5+1+6=12 columns
13*12=156 seats
.
Ed